# Hierarchical DLO Routing with Reinforcement Learning and In-Context Vision-language Models

### v1

@@ Abstract
Long-horizon routing tasks of deformable linear objects (DLOs), such as cables and ropes, are common in industrial assembly lines and everyday life. These tasks are particularly challenging because they require robots to manipulate DLO with long-horizon planning and reliable skill execution. Successfully completing such tasks demands adapting to their nonlinear dynamics, decomposing abstract routing goals, and generating multi-step plans composed of multiple skills, all of which require accurate high-level reasoning during execution. In this paper, we propose a fully autonomous hierarchical framework for solving challenging DLO routing tasks. Given an implicit or explicit routing goal expressed in language, our framework leverages vision-language models (VLMs) for in-context high-level reasoning to synthesize feasible plans, which are then executed by low-level skills trained via reinforcement learning. To improve robustness in long horizons, we further introduce a failure recovery mechanism that reorients the DLO into insertion-feasible states. Our approach generalizes to diverse scenes involving object attributes, spatial descriptions, as well as implicit language commands. It outperforms the next best baseline method by nearly 50% and achieves an overall success rate of 92.5% across long-horizon routing scenarios. Please refer to our project page: https://icra2026-dloroute.github.io/DLORoute/
@@ I Introduction
With the advent of VLMs, robotics has begun leveraging their powerful commonsense reasoning and scene understanding capabilities. These models can provide contextual knowledge, high-level task guidance, and even recovery strategies when a robot becomes stuck or encounters failure. For instance, VLM-PC [4] employs a VLM to guide a legged robot through unstructured environments by reasoning about progress and replanning dead-end cases, while relying on a locomotion controller for low-level actions. However, such approaches are unsuitable for DLO manipulation, where careless actions can damage the object—for example, pushing without regard to obstacles can lead to deformation or breakage. Thus, in addition to semantic reasoning for high-level planning, safe and reliable low-level control is essential for long-horizon DLO routing.
To address these challenges, we propose a hierarchical framework that integrates high-level planning via VLMs with reinforcement learning-based low-level control for DLO routing (Fig. 1). The objective is to route a DLO through multiple clips in a specified or natural order based on a language prompt. We design three core low-level skills: Insert and Pull for clip routing, and Flatten for failure recovery. Pull and Flatten are inherently safe as they move the DLO away from the environment and can be initialized from predefined motion primitives. In contrast, insertion requires precise navigation near a clip to maximize insertion success while avoiding collisions, so we train this skill using reinforcement learning to balance safety and accuracy.
@@
Figure 2: Pipeline of the proposed hierarchical DLO routing framework. The high-level VLM-based planner processes top-down scene images, task prompts, and auxiliary information to select appropriate skills, including routing (insertion and pulling) and failure recovery (flattening). Insertion is performed by a safe, low-level RL–based parameterized motion primitive for precise manipulation. A failure detection and recovery module (inference only) monitors execution to identify action, perception, or motion-planning failures and triggers replanning, enabling robust recovery and successful long-horizon routing. Note that side-view images shown in the second row are for visualization only; the planner receives only top-down views.
@@ III-A Problem Formulation
The planner is given an image of a scene $I_{scene}$ with three or more clips placed at different orientations and locations, along with a text description indicating the order in which it should route the DLO through. As shown in Fig. 2, the planner must generate a routing plan consistent with the provided description. The clip may have different attributes like color or spatial relation. The order will describe the color attributes, spatial relation, or an implicit order. We also consider a 4-clip setting to evaluate long-horizon generalization. In order to achieve the routing goal, a skill set containing possible skills is required. The planner should decide the next low-level skill to execute, the target clips to route through, and reason about task completion, given the scene image and history plan. Details about the planner are in Section III-C Additionally, the planner must decide when to stop execution by reasoning over task progress and comparing it with the goal specified in the initial instruction.
Figure 3: Illustration of the low-level action space for DLO routing skills. The primitive set includes Flatten (top left) and Pull actions (top right), and the Insert action (bottom). For insertion skill, the gripper (in orange) operates within a $0.16m\times 0.16m$ space with orientation ($\textbf{p}_{g}^{t}$, $\textbf{q}_{g}^{t}$) conditioned on the clip (in black) state ($\textbf{p}_{c}$, $\textbf{q}_{c}$). The DLO is represented by particles $\{\textbf{p}_{1},...,\textbf{p}_{n}\}$.
@@ III-B Training Low-level Skills
Low-level skills should robustly achieve their own manipulation goal. We designed three low-level skills available to the planners: Insert, Pull, Flatten, shown in Fig. 3. Each of them requires a specified target clip. Among these, insertion is the most crucial for cable routing, as it directly accomplishes the task objective. Successful insertion must adapt to the dynamics of deformable objects and be reactive during execution in the contact-sensitive regions. In contrast, pulling and flattening serve as auxiliary skills that adjust the DLO configuration to create more favorable conditions for insertion. Additionally, pulling and flattening move the DLO away from the clips, whereas insertion requires precise motion close to the clip and must achieve a high success rate to minimize repeated trials. Thus, for pulling and flattening, we employ a predefined motion primitive, while insertion is a parametrized motion primitive with its parameters optimized by a reinforcement learning agent trained in IsaacSim [12]. This allows insertion to leverage adaptive, learned behaviors for complex contact scenarios, while auxiliary skills remain simple and efficient.
We design the insertion motion primitive as shown in Fig. 3. First, the insertion primitive will grasp the corresponding sampled DLO segment $\mathbf{p}_{i}$ given a predicted grasping index $i$. After grasping, the robot follows a waypoint trajectory, where the via-points are defined by gripper poses $(\mathbf{q}_{g}^{t},\mathbf{p}_{g}^{t})$ at each time step $t$. The insertion RL policy learns to iteratively improve the insertion condition and can be executed for multiple timesteps, enabling closed-loop control and maximizing the insertion success.
We employ Soft Actor-Critic (SAC) [30] for training, with multi-layer perceptron (MLP) networks for both actor and critic networks. The reward function $r$ is defined as follows:
@@
Figure 4: Long-horizon routing examples in simulation using our VLM with failure reasoning and RL-trained low-level policy. (a) Implicit order forming a V-shape. (b–c) Fixed order based on color attributes (red, green, blue). (d) Fixed order determined by spatial reasoning. (e) Fixed order based on color attributes in a four-clip arrangement.
@@ III-C High-level In-context Learning
For a high-level VLM planner, we leverage in-context learning to adapt the robot to our task domain and output reliable low-level skill instructions. The planner takes as input a top-down scene image $I_{scene}$, and user-specified text about the routing order. In the first query, the planner outputs the predicted routing order (task plan). In subsequent queries, it outputs the predicted skills, target clips, and task progress information (skill-level execution). To achieve this, we instruct VLM to perform 1) Task progress reasoning, 2) low-level skills reasoning. Along with the user’s text, we also provide a scene description that includes the attributes of the DLO, clips, and environment to help improve scene understanding. Since the location of the DLO front segments is crucial for the routing with DLO head manipulation, we ask the planner to track the DLO head throughout the task and to recognize occlusions caused by clips when the DLO has already passed through them. This enables more accurate reasoning about task progress, target clip selection, and completion status. To further enhance decision making, we employ the chain-of-thought (CoT) prompting [31] to ask the planner to output the zoom-in view analysis, thereby determining DLO–clip interaction before predicting the next skill or routing order.
Skill reasoning helps planners distinguish between available skills and make precise selections. An elaborated definition of each skill, along with verbal usage examples is given to VLM. Additionally, we provide two example images corresponding to insertion and pulling, illustrating the detailed scenario when these skills should be invoked.
@@ III-D Hierarchical Planning & Failure Recovery
During the DLO routing rollout with only insertion and pulling skills, we observed that when pulling the DLO toward the next clip, the DLO head inevitably twisted or drifted far away from the clip opening. Under such challenging DLO states, insertion attempts will consistently fail. To address this problem, we introduce a failure recovery mechanism. The recovery module consists of a low-level rescue skill, Flatten, a simple yet effective failure recovery skill that reintializes the DLO to an insertion-possible state. In addition to providing the definition and usage example of flattening skill, we set a counter of consecutive insertion attempts and feed this status summary to the planner. We then impose a step limit to prevent getting stuck by infinitely executing the same skill and incorporating this constraint into the CoT prompting. Failure detection and recovery are automatically predicted by the planner from observations and history, without human intervention. With this design, even if the DLO head aligns nearly parallel with the clip’s long axis, the high-level planner will not get stuck and can invoke recovery, reorient the DLO, and resume insertion successfully.
@@ IV Experiments
We perform the experiments to answer the following research questions: Q1: Can our approach with a hierarchical framework excel at long-horizon deformable manipulation tasks? Q2: Does failure reasoning boost the performance during long-horizon reasoning? Q3: Would our approach generalize to unseen scenarios (e.g., more clips)? Q4: Does our approach outperform baselines using a visual baseline as the low-level policy? In this paper, we train our low-level RL policy in IsaacSim [12] and GarmentLab [32] and use GPT-5 [33] with low reasoning effort as the VLMs planner for our experiments.
@@ IV-A1 Low-level Skills Setup
Metrics Visual Baseline RL Policy (Ours)
@@
TABLE I: Real-world evaluation on open-loop methods under different friction environments.
High Level Low Level Eval. Metric Implicit Fixed (Spatial) Fixed (Attribute) 4-Clip
VLM w/ Failure Reasoning Visual Baseline Success Rate (%) $\uparrow$ 20 10 0 10
Avg. #Clips Ins. $\uparrow$ 1.7 0.7 1.1 1.3
Avg. Epi. Length 9.6 7.6 7.4 9.2
VLM RL policy Success Rate (%) $\uparrow$ 70 10 10 50
Avg. #Clips Ins. $\uparrow$ 2.5 1.5 1.3 3.1
Avg. Epi. Length 14.1 21.5 20.4 27.7
Fixed Order RL Policy Success Rate (%) $\uparrow$ 80 20 20 60
Avg. #Clips Ins. $\uparrow$ 2.6 1.6 1.5 3.2
Avg. Epi. Length 10.1 9.6 8.9 13.7
VLM w/ Failure Reasoning RL policy Success Rate (%) $\uparrow$ 80 90 100 100
Avg. #Clips Ins. $\uparrow$ 2.8 2.9 3 4
Avg. Epi. Length 12.8 18.4 17.5 17.9
TABLE II: Simulation result for long-horizon DLO routing. Evaluation metrics include success rate (higher is better), average number of clips inserted (higher is better), and average episode length. Our hierarchical framework with failure reasoning achieves the best performance across settings.
@@ IV-A2 Evaluation Details
We evaluate the low-level policies using 100 different scenes of randomized DLO poses and clips with the original scale. Average distance (avg. dis.) is the average signed endpoint distance $\pm d_{\mathit{ceil}}$. It is positive when $\mathit{rope\_out}$ and negative otherwise. The success is recorded when the signed endpoint distance $\pm d_{\mathit{ceil}}$ is larger than +2cm after the motion primitive finishes.
As for long-horizon DLO routing, we evaluate our baselines with four different DLO routing strategies: implicit orders, fixed orders with spatial descriptions or color attributes, and fixed orders with color attributes under 4-clip settings, as shown in Fig. 4. Note that the 4-clip setting is created by adding one more clip to the front or back of the clip arrangement in the implicit routing order setting. The DLO must fully pass through each clip, and the DLO head must emerge from the other side. A success case is recorded when the DLO is routed through all clips in the scene with the correct order. Average episode length (Avg. Epi. Length) means the number of pick-and-place steps executed. Average number of clips inserted (Avg. #Clips Ins.) shows the average number of clips inserted when the episode terminates. An episode may be terminated and be considered a failure case when 1) a collision is detected, 2) the maximum timeout has been reached, or 3) early termination with some clips remaining not inserted.
@@ IV-B Baselines
  • Visual baseline (VB): A heuristic approach that imitates a human expert’s insertion strategy. We design a straightforward insertion action for different DLO shapes and clip sizes, where we first place the DLO in front of the clip opening and insert along the center line of the clip.
For long-horizon DLO routing evaluation, we employ several heuristic methods and conduct ablation studies under various challenging clip arrangements.
@@
  • VB as low-level policy: substitutes the low-level RL policy with the heuristic strategy that mimics human insertion but is less considerate of the environment, allowing us to evaluate performance under a simpler motion primitive.
@@ IV-C Simulation Experiments
Q1: Can our approach with a hierarchical framework excel at long-horizon deformable manipulation tasks? The hierarchical planning has greatly improved the performance compared with the fixed order heuristic baseline. As shown in Table II, our method excels for most tasks, and shows a better success rate and higher number of clips inserted across all evaluation settings, particularly under challenging conditions such as Fixed Spatial, Fixed Attribute, and 4-Clip. While our method requires longer episode lengths, the improved robustness and task completion demonstrate that the hierarchical framework excels at long-horizon deformable manipulation tasks. The fixed order baseline achieves comparative performance in an implicit setting, primarily because insertion in this scenario is less challenging. In order to create cases for natural ordering, the clip opening angles within this experiment set are similar to each other to ensure mild turns between the clips. Such settings reduce the need for failure recovery, as the fixed order works well already. However, when it comes to challenging cases with sharper turns, the insertion needs more trials or re-initialization for the DLO to get through. Certain failure cases for ours are shown in Fig. 6, the high-level planner issues an early termination before the DLO has been fully inserted into the last clip.
@@
Q2: Incorporating failure reasoning boosts our performance over the plain VLM policy. Across all evaluation settings, our method achieves higher success rates and more clips inserted, with the most substantial gains in challenging long-horizon scenarios such as Fixed Spatial, Fixed Attribute, and 4-Clip. Moreover, episodes are shorter in these cases, indicating that the agent recovers more efficiently from errors. In contrast, an agent without failure reasoning can only try several times under insertion-challenging scenarios and keep getting failure results. Additionally, scenarios such as Fixed Spatial and Fixed Attribute are relieved from the mild turn preset condition and contain over 90-degree turns between clips. This makes insertion much harder without recovering from failure. These results demonstrate that failure reasoning is a key factor in boosting long-horizon reasoning performance.
@@
Q3: Our approach can extend beyond the current clip setting to a more challenging 4-clip setting. Our hierarchical framework with failure reasoning achieves a perfect success rate and consistently inserts all clips, demonstrating strong generalization to longer-horizon tasks. By contrast, both the VLM-only policy and the Fixed Order baseline suffer notable drops in success rate and efficiency. These results indicate that our approach scales robustly to scenarios involving more clips. Although this setting is extended from relatively simple routing scenarios, the added clip usually forms a large angle compared with the previous clip, making the setting much harder for simple insertion strategies.
@@
Q4: Comparing against the visual baseline shows that the low-level policy plays a critical role in long-horizon deformable manipulation. While the heuristic insertion baseline quickly fails due to its lack of environmental awareness, the RL low-level policy adapts actions to task progress and scene dynamics, enabling precise and robust insertions. This distinction becomes most evident in the 4-Clip setting, where the visual baseline achieves only 10% success while our approach reaches 100%, demonstrating that task-aware low-level skill is essential for scaling to complex scenarios.

### v2

@@ Abstract
Long-horizon routing tasks of deformable linear objects (DLOs), such as cables and ropes, are common in industrial assembly lines and everyday life. These tasks are particularly challenging because they require robots to manipulate DLO with long-horizon planning and reliable skill execution. Successfully completing such tasks demands adapting to their nonlinear dynamics, decomposing abstract routing goals, and generating multi-step plans composed of multiple skills, all of which require accurate high-level reasoning during execution. In this paper, we propose a fully autonomous hierarchical framework for solving challenging DLO routing tasks. Given an implicit or explicit routing goal expressed in language, our framework leverages vision-language models (VLMs) for in-context high-level reasoning to synthesize feasible plans, which are then executed by low-level skills trained via reinforcement learning. To improve robustness over long horizons, we further introduce a failure recovery mechanism that reorients the DLO into insertion-feasible states. Our approach generalizes to diverse scenes involving object attributes, spatial descriptions, implicit language commands, and extended 5-clip settings. It achieves an overall success rate of 92% across long-horizon routing scenarios. Please refer to our project page: https://icra2026-dloroute.github.io/DLORoute/
@@ I Introduction
With the advent of VLMs, robotics has begun leveraging their powerful commonsense reasoning and scene understanding capabilities. These models can provide contextual knowledge, high-level task guidance, and even recovery strategies when a robot becomes stuck or encounters failure. For instance, VLM-PC [2] employs a VLM to guide a legged robot through unstructured environments by reasoning about progress and replanning dead-end cases, while relying on a locomotion controller for low-level actions. However, such approaches are unsuitable for DLO manipulation, where careless actions can damage the object. For example, pushing without regard to obstacles can lead to deformation or breakage. Thus, in addition to semantic reasoning for high-level planning, safe and reliable low-level control is essential for long-horizon DLO routing.
To address these challenges, we propose a hierarchical framework that integrates high-level planning via VLMs with reinforcement learning-based low-level control for DLO routing (Fig. 1). The objective is to route a DLO through multiple clips in a specified or natural order based on a language prompt. Our framework is designed for longer-horizon tasks and generalizes from 3-clip to more challenging 4- and 5-clip settings. We design three core low-level skills: Insert and Pull for clip routing, and Flatten for failure recovery. Pull and Flatten are inherently safe as they move the DLO away from the environment and can be initialized from predefined motion primitives. In contrast, insertion requires precise navigation near a clip to maximize insertion success while avoiding collisions, so we train this skill using reinforcement learning to balance safety and accuracy.
@@
Figure 2: Pipeline of the proposed hierarchical DLO routing framework. The high-level VLM-based planner processes top-down scene images, task prompts, and auxiliary information to select appropriate skills, including routing (insertion and pulling) and failure recovery (flattening). Insertion is performed by a safe, low-level RL–based parameterized motion primitive for precise manipulation. A failure detection and recovery module (inference only) monitors execution to identify failures and triggers replanning, enabling robust recovery and successful long-horizon routing. Note that the planner receives only top-down views.
@@ III-A Problem Formulation
The planner is given an image of a scene $I_{scene}$ with three or more clips placed at different orientations and locations, along with a text description indicating the order in which it should route the DLO through. As shown in Fig. 2, the planner must generate a routing plan consistent with the provided description. The clip may have different attributes like color or spatial relation. The order will describe the color attributes, spatial relation, or an implicit order. We also consider extended multi-clip settings (up to 5 clips) to evaluate long-horizon generalization. In order to achieve the routing goal, a skill set containing possible skills is required. The planner should decide the next low-level skill to execute, the target clips to route through, and reason about task completion, given the scene image and history plan. Details about the planner are in Section III-C Additionally, the planner must decide when to stop execution by reasoning over task progress and comparing it with the goal specified in the initial instruction.
Figure 3: Illustration of the low-level action space for DLO routing skills. The primitive set includes Flatten (top left) and Pull actions (top right), and the Insert action (bottom). For insertion skill, the gripper (in red) operates within a $0.16m\times 0.16m$ space with orientation ($\textbf{p}_{g}^{t}$, $\textbf{q}_{g}^{t}$) conditioned on the clip (in black) state ($\textbf{p}_{c}$, $\textbf{q}_{c}$). The DLO is represented by particles $\{\textbf{p}_{1},...,\textbf{p}_{n}\}$.
@@ III-B Training Low-level Skills
Low-level skills should robustly achieve their own manipulation goal. We designed three low-level skills available to the planners: Insert, Pull, Flatten, shown in Fig. 3. Each of them requires a specified target clip. Among these, insertion is the most crucial for DLO routing, as it directly accomplishes the task objective. Successful insertion must adapt to the dynamics of deformable objects and be reactive during execution in the contact-sensitive regions. In contrast, pulling and flattening serve as auxiliary skills that adjust the DLO configuration to create more favorable conditions for insertion. Additionally, pulling and flattening move the DLO away from the clips, whereas insertion requires precise motion close to the clip and must achieve a high success rate to minimize repeated trials. Thus, for pulling and flattening, we employ a predefined motion primitive, while insertion is a parametrized motion primitive with its parameters optimized by a reinforcement learning agent trained in IsaacSim [23]. This allows insertion to leverage adaptive, learned behaviors for complex contact scenarios, while auxiliary skills remain simple and efficient.
We design the insertion motion primitive as shown in Fig. 3. First, the insertion primitive will grasp the corresponding sampled DLO segment $\mathbf{p}_{i}$ given a predicted grasping index $i$. After grasping, the robot follows a waypoint trajectory, with via points defined by gripper poses $(\mathbf{q}_{g}^{t},\mathbf{p}_{g}^{t})$ at each time step $t$. The RL policy observation consists of particle-based DLO state $\mathbf{p}_{1:n}$, the clip pose and scale, and a binary indicator rope_inrope\_in indicating partial DLO insertion. The RL policy learns to iteratively improve the insertion condition and can be executed for multiple timesteps, enabling closed-loop control and maximizing the insertion success.
We employ Soft Actor-Critic (SAC) [9] for training, with multi-layer perceptron (MLP) networks for both the actor and the critic. The reward function $r$ is defined as follows:
@@ III-C High-level In-context Learning
For a high-level VLM planner, we leverage in-context learning to adapt the robot to our task domain and output reliable low-level skill instructions. The VLM planner operates strictly at the high-level layer. It decides the routing order and selects among a predefined set of skills and target clips, but does not generate low-level actions. To predict the routing order, the planner takes as input a top-down scene image $I_{scene}$ and a user-specified routing instruction. In subsequent queries, it outputs a structured decision: selected skill, target clip, and task completion flag. To achieve this, we instruct VLM to perform 1) Task progress reasoning, 2) low-level skills reasoning. Along with the user’s text, we provide a scene description that includes the attributes of the DLO, clips, and environment to help improve scene understanding. Since the locations of the DLO front segments are crucial for the routing with DLO head manipulation, we ask the planner to track the DLO head throughout the task and to recognize occlusions caused by clips when the DLO has already passed through them. This enables more accurate reasoning about task progress, target clip selection, and completion status. To further enhance decision making, we employ chain-of-thought (CoT) prompting [34] to ask the planner to analyze the zoom-in view, thereby determining DLO–clip interaction before predicting the next skill and target clip. Skill reasoning helps the planner distinguish between available skills and make precise selections. We provide the VLM with structured definitions for each skill (Insert, Pull, Flatten), including their functional objectives, usage conditions, and counter-examples specifying when they should not be invoked. For instance, insertion is conditioned on the rope head being close to and roughly aligned with the clip opening. We provide two in-context examples for insertion and pulling scenarios, each consisting of an image and its corresponding skill, similar to the leftmost examples in Figure 5. The prompt also specifies criteria for insertion success and task success.
@@ III-D Hierarchical Planning & Failure Recovery
During the DLO routing rollout with only insertion and pulling skills, we observed that when pulling the DLO toward the next clip, the DLO head may twist or drift far away from the clip opening. Under such states, insertion attempts will consistently fail. To address this problem, we introduce a failure recovery mechanism. The recovery module consists of a low-level rescue skill, Flatten, a simple yet effective failure-recovery skill that reintializes the DLO to an insertion-possible state. In addition to providing the definition and usage example of flattening skill, we set a counter of consecutive insertion attempts and feed this status to the planner. We then impose a step limit to prevent getting stuck by infinitely executing the same skill and incorporating this constraint into the CoT prompting. Failure detection and recovery are automatically predicted by the planner from observations and history, without human intervention. With this design, even if the DLO head aligns with the clip’s long axis, making insertion difficult, the planner can invoke recovery to reorient the DLO and resume insertion.
Figure 4: Long-horizon routing examples in simulation using our VLM with failure reasoning and RL policy. (a) Implicit order forming a V-shape. (b–c) Fixed order based on color attributes (red, green, blue). (d) Fixed order determined by spatial reasoning. (e) Fixed order based on color attributes in a 5-clip arrangement.
@@ IV Experiments
We perform the experiments to answer the following research questions: Q1: Can our approach with a hierarchical framework excel at long-horizon deformable manipulation tasks? Q2: Does failure reasoning boost the performance during long-horizon reasoning? Q3: Would our approach generalize to unseen scenarios (e.g., more clips)? Q4: Does our approach outperform baselines using a heuristic policy as the low-level policy? In this paper, we train our low-level RL policy in IsaacSim [23] and GarmentLab [20] and use GPT-5 [24] with low reasoning effort as the VLMs planner for our experiments.
@@ IV-A1 Low-level Skills Setup
Metrics Heuristic Policy RL Policy (Ours)
@@
TABLE I: Simulation result for rope insertion primitive.
High Level Low Level Implicit Fixed (Spatial) Fixed (Attr.) 4/5-Clip
SR $\uparrow$ Clips $\uparrow$ Epi SR $\uparrow$ Clips $\uparrow$ Epi SR $\uparrow$ Clips $\uparrow$ Epi SR $\uparrow$ Clips $\uparrow$ Epi
VLM w/ Failure Reasoning Heuristic Policy 13 1.3 8.0 7 0.7 6.9 0 0.9 7.3 7 1.3 9.1
VLM w/o Failure Reasoning RL Policy 47 2.0 16.2 7 1.3 22.7 7 1.2 23.3 47 3.1 28.4
Fixed Order RL Policy 53 2.1 9.7 13 1.3 8.7 13 1.3 8.9 60 3.4 14.6
Symbolic Planner RL Policy 68 2.3 10.0 53 2.1 11.5 53 2.1 12.0 100 4.3 21.5
VLM w/ Failure Reasoning RL Policy 80 2.7 13.9 93 2.9 18.3 93 2.9 17.7 100 4.3 20.7
TABLE II: Simulation results for long-horizon DLO routing. SR = success rate (%), Clips = average number of clips inserted, Epi = average episode length. Higher SR and Clips indicate better performance.
@@ IV-A2 Evaluation Details
As shown in Tab. I, we evaluate the low-level policies on 100 different scenes with randomized DLO poses and clips at the original scale, matching the real-world clips used in our experiments. Average distance (avg. dis.) is the average signed endpoint distance $\pm d_{\mathit{ceil}}$. It is positive when $\mathit{rope\_out}$ and negative otherwise. The success is recorded when the signed endpoint distance $\pm d_{\mathit{ceil}}$ is larger than +2cm after the motion primitive finishes.
As for long-horizon DLO routing, we evaluate our baselines with four different DLO routing strategies: implicit orders, fixed orders with spatial descriptions or color attributes, and fixed orders with color attributes under 4/5-clip settings, as shown in Fig. 4. Each strategy is evaluated over 15 trials with different clip layouts. The DLO must fully pass through each clip, and the DLO head must emerge from the other side. A success case is recorded when the DLO is routed through all clips in the scene with the correct order. Average episode length means the number of pick-and-place steps executed. Average number of clips inserted shows the average number of clips inserted when the episode terminates. An episode may be terminated and be considered a failure case when 1) a collision is detected, 2) the maximum timeout has been reached, or 3) early termination with some clips remaining not fully inserted.
@@ IV-B Baselines
  • Heuristic Policy: An insertion method that imitates a human expert’s strategy. The robot grasps the DLO head and places it in front of the clip, then regrasps further along the DLO and inserts it along the clip centerline.
For long-horizon DLO routing evaluation, we employ several heuristic methods and conduct ablation studies.
@@
  • Heuristic policy as low-level policy: substitutes the low-level RL policy with the heuristic strategy which is less considerate of the environment, allowing us to evaluate performance under a simpler motion primitive.
  • Symbolic Planner: The planner uses ground-truth clip poses and the rope head position to decide whether to execute Insert or Pull. The routing order is determined using a greedy strategy that selects the next closest clip. Insertion is considered successful when the rope head passes through the clip opening; otherwise, it retries insertion and triggers Flatten after three failures.
@@ IV-C Simulation Experiments
Q1: The hierarchical planning has greatly improved the performance compared with the fixed order and heuristic policy baselines. As shown in Table II, our method excels across all tasks, achieving a higher success rate and a greater number of clips inserted across all evaluation settings, even under challenging conditions. While our method requires longer episode lengths, the improved robustness and task completion demonstrate that the hierarchical framework excels at long-horizon deformable manipulation tasks. The symbolic planner baseline highlights the limitations of purely heuristic planning. Its greedy routing strategy cannot process language instructions and may select the closest clip instead of the correct target, particularly under challenging conditions such as Implicit, Fixed Spatial, Fixed Attribute. When the greedy order happens to match the correct sequence (e.g., in the 4/5-clip setting), it can achieve high success rates but still results in longer episodes because it repeatedly attempts insertion before recovery. In contrast, our method detects failures earlier and invokes the Flatten action more efficiently. Certain failure cases for ours are shown in Fig. 6, the high-level planner issues an early termination before the DLO has been fully inserted into the last clip.
@@
Q2: Incorporating failure reasoning boosts our performance over the plain VLM policy. Across all evaluation settings, our method achieves higher success rates and more clips inserted, with the most substantial gains in all challenging long-horizon scenarios. Moreover, episodes are shorter in these cases, indicating that the agent recovers more efficiently from errors. In contrast, an agent without failure reasoning can only try several times under insertion-challenging scenarios and keep getting failure results. As shown in Table II, baselines without failure reasoning (e.g., the plain VLM policy and Fixed Order) perform reasonably well in the implicit setting. In this case, the clip opening angles are similar, producing mild turns between clips and forming a natural routing order. Such configurations reduce the need for failure recovery, allowing these baselines to succeed. However, scenarios such as Fixed Spatial and Fixed Attribute are relieved from the mild turn preset condition and contain over 90-degree turns between clips. This makes insertion much harder without recovering from failure. These results demonstrate that failure reasoning is a key factor in boosting long-horizon reasoning performance.
@@
Q3: Our approach can extend beyond the current clip setting to a more challenging 4/5-clip settings. Although this setting is extended from relatively simple routing scenarios, the added clip usually forms a large angle compared with the previous clip, making the setting much harder for simple insertion strategies. Despite this difficulty, our hierarchical framework with failure reasoning achieves a perfect success rate and consistently inserts all clips, demonstrating strong generalization to longer-horizon tasks. By contrast, both the VLM-only policy and the Fixed Order baseline suffer notable drops in success rate and efficiency. These results indicate that our approach scales well with the number of clips, as the high-level planner reliably predicts next skills while the low-level policy performs consistently across different clips.
@@
Q4: Comparing against the heuristic policy shows that the low-level policy plays a critical role in long-horizon deformable manipulation. While the heuristic insertion baseline quickly fails due to its lack of environmental awareness, the RL low-level policy adapts actions to task progress and scene dynamics, enabling precise and robust insertions. This distinction becomes most evident in the 4/5-Clip setting, where the heuristic policy achieves only 7% success while our approach reaches 100%, demonstrating that task-aware low-level skill is essential for scaling to complex scenarios.